# Daria Semochkina School of Mathematical Sciences

### v1

@@ 4 Example: SIS model
The approach adopted here is to add additional information about the parameters in the form of prior information. This prior information is specified by using information from other sources (e.g. from expert knowledge or previous experience of these parameters). Note that while the term ‘prior’ may in ordinary usage be associated with information available before collecting data (i.e. targets), in the Bayesian paradigm it is the fact that information other than contained in the targets may be incorporated that is important.
In this case, an informative prior is used. We started with using uniform priors for each parameter. However, only one posterior distribution (for parameter p) was distinct from the priors. The next step was to impose normal prior for two parameter $c\sim N(9,1)\text{ and }p\sim N(0.07,0.01)$ (See rhs of Figure 2 for results). The chains were compared using the Gelman and Rubin’s convergence diagnostic[5]. The no prior chains scored a point estimate 2.58 with an upper confidence interval of 5.05 and the chains with normal priors scored a 1.05 point estimate with upped C.I. of 1.15. The posterior distributions for the parameters were very close to the imposed priors, highlighting the fact that there is no identifying information in the model itself.
@@
In summary, in this section we have illustrated how the problem of non-identifiability can arise in even a relatively simple model, and shown how the use of additional information solves this problem. It was after initial runs that the modeller will detect that non-identifiability is occurring, and by carefully considering what information to add regarding other parameters, identifiability may be achieved. In practice, such additional information may come from, for example, monitoring a number of individual cases of the disease to better quantify the time to recovery.
@@ 5 Example: HPV model
The second model that we use in this paper to illustrate the challenges of the Bayesian MCMC approach to model calibration is an individual-based model, also called an agent-based model. An individual-based model simulates the characteristics of each individual and keeps track of each individual’s history, recording all information relevant to the disease in question.
The advantages of agent based models are that they can incorporate heterogeneity (through each agent), they can be physically realistic, and they allow for interventions that impact on individuals differently. They can commonly be used when modelling dynamic aspects of infectious diseases as described for example in [6].
The challenges associated with these models is that they take more computer time to run, they can be difficult to build, and of interest to this example they can be harder to calibrate.
The example we use here is based on the human papillomavirus (HPV) which causes cervical cancer. We follow the model developed by Kim et al. [7] which considers transitions between Healthy state, HPV infection, pre-malignant changes called cervical intra-epithelial neoplasia (CIN) 1, CIN 2 and CIN 3 (which are different levels of severity) and, finally, cancer (see Figure 3 for details).
The model allows both forward and backward transitions which leads to a complex structure. Additionally, the model’s parameters vary by strain (i.e. transition probabilities to CIN states depend on the strain of HPV infection), and therefore the number of parameters for calibration is high. We have deliberately selected this HPV model for this paper because it is likely to exhibit many of the challenges that were discussed in Section 2.1.
@@
5.1 Data
@@
We assume that all of the targets have normal distributions $f_{k}\sim N(d_{k},\sigma_{k}),$ centred at the observed data $d_{k}$ and with variances $\sigma_{k}$ that were calculated from the data. We assume all targets to be independent. Because of that assumption, our likelihood is just a product of individual likelihoods
@@
5.2 Calibration using Metropolis-Hastings
In this section, we present some results of applying the Metropolis-Hastings algorithm to the HPV model calibration. They illustrate some of the problems associated with calibrating complex models using the Metropolis-Hastings algorithm described in Section 2.1.
The trade-off between adequate mixing and good acceptance rate is a significant source of difficulty. We have considered many different configurations of the algorithm in the process of calibrating the HPV model. We present three choices in the remainder of this section.
5.2.1 The Metropolis-Hastings algorithm with independent normal proposals
The first attempt at calibration was made where all parameters were subject to random changes on each step of the algorithm (i.e. $s=0$ case in Appendix Algorithm 1). This, however, is a less than ideal approach. As was illustrated in Figure 1(a) of Section 2, when you have two competing transition probabilities, successfully calibrating both can be impossible if the data available for calibration is population-averaged. Unfortunately, in our case, the only data available were the distributions of population-level observations. This is a common problem of incompatibility of population-level data (the most commonly collected and available) with agent-based models. As a necessary restriction, all backward transition probabilities were constrained at high-likelihood values for all subsequent realisations of the algorithm.
To further address the problem of low acceptance rate, we utilise the $s$ parameter in Appendix Algorithm 1. We will refer to this case as the ‘partial independent normal proposals’. At each step of the algorithm, we select $s$ parameters to update, while the remaining $n-s$ parameters are set to their values at the previous step. It is intuitively obvious that the algorithm is more likely to make a reasonable proposal (at random) for fewer parameters, comparing to the full parameter space of $n$ parameters. The number of parameters proposed to be updated was the same at each step. The main benefit of this approach was to increase the acceptance rate to $\sim$15%, compared to $<$1%.
5.2.2 The Metropolis-Hastings algorithm with informative priors
To further improve the performance of Appendix Algorithm 1 and particularly target the problem of the non-identifiability of the parameters, we introduce informative priors. There are many ways to incorporate any information available at a given time into priors. If the priors for each parameter are assumed to be independent, then the prior $\pi(\boldsymbol{x})$ in Appendix Algorithm 1 is just a product of one-dimensional priors $\pi(\boldsymbol{x})=\prod_{i=1}^{n}\pi_{i}(x_{i}),$ where each $\pi_{i}$ is a prior distribution for $i$’s parameter. The choice of priors represents our beliefs about parameters without taking new evidence into account. Gamma distributions were chosen to be fitted to the parameters because of their $[0,\infty)$ support and their shape. Priors on the model’s parameters were imposed to limit the algorithm’s search to (what we know to be) a plausible region of the space. The uniqueness of such a region can depend on many things. Possible sources of information from which one could construct priors also include expert opinions and alternative data sets.
Having informative priors produced the best results for the HPV model calibration. We have chosen to present only those results in detail in the next section, focusing on demonstrating the effects of having an informative prior on a complex parameter space with correlated parameters.
6 Algorithm results with informative priors
As highlighted above, the prior distribution can be thought of as ‘other information’ about the parameters than is provided by the likelihood. In this case, structural priors can be placed upon the parameter space by observing that large sections are not compatible with plausible data values. As highlighted in Section 5.2.1, plausible values had also to be chosen for the backward transitions.
To improve the convergence of chains and address the problems described in Section 2.1, structural priors on the joint parameter space were imposed. The chains with the best GOF were chosen out of all the chains obtained using all previous attempts described in Section 5.2. Those chains, containing the information on plausible parameter values, were used to fit the priors. The gamma distributions were fitted so that $x_{i}\sim\Gamma(\widehat{\alpha}_{i},\widehat{\beta}_{i})$, where $x_{i}$ is the $i^{th}$ parameter and $\widehat{\alpha}_{i}$ and $\widehat{\beta}_{i}$ were estimated from the identified chain.
The resulting chains of the Metropolis-Hastings algorithm with informative priors and partial independent normal proposals for four selected parameters can be found in Figure 4 along with their densities. Although here we focus on four of the most interesting parameters, the resulting densities for all 26 parameters are presented in the Appendix Figure 1. It is visible from the density plots in Figure 4 that some multi-modality exists in the space. All three chains are recognising the multi-modality and sampling from the same distribution. This multi-modality was not picked up by the original chain, chosen to inform the gamma priors. This highlights the importance of prior elicitation in complex models. The balance between informing some parameters from the data or informing them through the priors could be drastically shifted towards the priors in a situation of non-identifiability[4] or just general lack of information in the available data.
Figure 4: Results of running the Metropolis-Hastings algorithm with informative priors and partial independent normal proposals for four selected parameters. For each parameter, both the chains and distributions are presented. Each parameter represents a multiplier to a baseline monthly probability of transitioning from one state to another, conditional on a type of infection (see Figure 3 for details). The conditioning on the infection type is presented in parentheses. See Figure 3 and Kim et al. [7] for details on baseline probabilities of transitioning.
Despite the complications associated with the application of the Metropolis-Hastings algorithm to the HPV model calibration, it is noted that when compared to the original method of Kim et al. [7] the GOF was consistently better for all chains (low scores being better), staying near 20, fluctuating up and down around that value. Despite the problems, this was true for all approaches, even the ones that we have chosen not to present here. If we compare this to the simple Monte Carlo approach of Kim et al. [7], the best fitting parameter values had GOF values around 122 for the Irish data (with minimal value at 72 and maximal value at 185 in a sample of 133 best fitting values). More details can be found in the Appendix.
@@ 6 Algorithm results with informative priors
As was mentioned in Section 5.1, the likelihood is a product of independent normal density functions and GOF is the scaled likelihood on the log scale. The details of the means and standard deviations can be found in Appendix Table 1. If we simplify the GOF measure, we can see that it is $\chi^{2}$ with 31 degrees of freedom (assuming independence).
@@
The corresponding p-values are 93.58% and $\ll 0.01\%$ for our approach (GOF=20) and Kim et al. [7] approach (GOF=122) respectively. However, the fit of our approach appears to be almost ‘too good’ relative to $\chi^{2}_{31}$. That can be explained by the correlation in the targets, so if one is fitting well, a few others are also fitting well. If we remove the assumption of independence, that measure of GOF would have a different distribution (depending on the correlations), but the mean of that distribution will be shifted towards zero compared to $E(\chi^{2}_{k=31})=31$. This will make the value of GOF=20 more realistic.
In general, the parameter values identified by the original approach were less concentrated, compared to posterior distributions produced by the Metropolis-Hastings algorithm. Most of the ‘best fitting’ values spanned the entire search range and didn’t identify areas with higher likelihood. Additionally, some of the good-fitting values identified by the Metropolis-Hastings algorithm were outside of the originally specified ranges which was also missed by the original approach. Overall, the Metropolis-Hastings algorithm was able to find better-fitting parameter combinations, the distributions identified had smaller variances and the search required fewer model runs to produce those results.
7 Discussion
Given the advantages of a Bayesian stochastic perspective, and in particular here an MCMC approach, the fact that formal convergence is not always easily obtained without substantial additional work requires much attention. Such work includes better data collection or data representation in terms of priors as well as tuning the MCMC to the problem.
The Bayesian approach to model calibration is becoming more popular[8] and more research is focusing on MCMC techniques in particular. Several calibration approaches become more difficult to apply in the context of stochastic models, such as stochastic microsimulations or agent-based models[8].
One of the main observations in this work was that it is possible to explicitly incorporate additional information into the inference in the form of priors during the calibration process and that this had a significant effect on the algorithm’s performance. This has a big impact on the issue of non-identifiability. Once the additional information was included (here through the revision of the prior), the algorithm was not impaired by the presence of some local multi-modality and all three chains sampled from the same distribution. This led to a desirable effect of convergence that was coupled with the improved acceptance rate of the algorithm and visible lack of autocorrelation in the sample.
The fact that the application of the Metropolis-Hastings algorithm yielded results with a lower GOF than the simple Monte Carlo approach is not surprising, given that the Metropolis-Hastings algorithm is designed to lead the sampler to a high likelihood area, whether it is a global or a local maximum. Additionally, this technique provides probabilistic estimates of the parameters of interest and reflects their joint uncertainty. Usually, the acceptance criterion is applied after the fact, which leads to inefficiency.
In a population-level model of Hepatitis C infection a similar process of calibration was performed [9]. This case had none of the problems that we had with the HPV model calibration. This was mostly because good priors were provided by an expert. Furthermore, because the Hepatitis C model did not have any backward transitions, there were fewer problems with identifiability, compared to the HPV model. Another possible advantage of that model was its relative simplicity, making it difficult to over-parameterise, compared to the individual-based models. This allowed us to try multiple strategies, before selecting the best one. We hypothesise that the relatively simpler structure of the Hepatitis C model, compared with the HPV model, yields a smoother probabilistic link between the model parameters and outputs. Of course, model choice is very influential in model calibration and different model types might yield different calibration results.
Funding
Acknowledgements
We would like to thank MACSI (University of Limerick) and the Health Research Board who have provided the funding for this research. We are also grateful to Jane Kim for sharing the code from their paper[7]. Daria Semochkina would also like to thank Paul Smith for his advice on article preparation and many helpful discussions.
Appendix A Appendices
A.1 Metropolis-Hastings algorithm
The central idea of the Metro[polius-Hastings algorithm is to update locally, i.e. use the accepted parameter value $\boldsymbol{x}_{(t)}$ at time step $t$ to generate the next proposed value. This makes it easier to think of a suitable proposal step (that is, of course, conditional), albeit with the disadvantage of producing a Markov chain as opposed to an independent realisation. This Markov chain converges to the posterior distribution of interest after enough steps. The sample values are taken from the Markov chain after convergence. Below we present the M-H algorithm we used to calibrate the HPV model.
Algorithm 1 HPV model calibration using Metropolis-Hastings.
Starting from $\boldsymbol{x}^{(0)}=(x^{(0)}_{1},x^{(0)}_{2},...,x^{(0)}_{n})$ for $t=1$ to $N=N_{steps}$ do
1:Draw $\boldsymbol{x}\sim q(\cdot|\boldsymbol{x}^{(t-1)}),$ where $q$ is the proposal distribution
2:For some $s<n$ randomly select $n-s$ parameters $x_{j_{1}},...x_{j_{n-s}}$ and set them to previous values $x_{j_{1}}=x^{(t-1)}_{j_{1}},...,x_{j_{n-s}}=x^{(t-1)}_{j_{n-s}}$
3:Run the model for a cohort of 25,000 women, given the new parameter set $\boldsymbol{x}$
4:Calculate the likelihood of the newly proposed step $f(\boldsymbol{x})$
5:Compute
$p(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})=\min\left\{1,\frac{\pi(\boldsymbol{x})f(\boldsymbol{x})q(\boldsymbol{x}^{(t-1)}|\boldsymbol{x})}{\pi(\boldsymbol{x}^{(t-1)})f(\boldsymbol{x}^{(t-1)})q(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})}\right\},$ (1)
where $\pi(\boldsymbol{x})$ is the prior distribution
6:With probability $p(\boldsymbol{x}|\boldsymbol{x}^{(t-1)})$ accept the step and set $\boldsymbol{x}^{(t)}=\boldsymbol{x}$, otherwise reject the step and let $\boldsymbol{x}^{(t)}=\boldsymbol{x}^{(t-1)}$
End
A.2 HPV model
In this appendix, we present further details about the HPV model, described in Section 5. In Appendix Table 2 we present the best five parameter combination that produced the lowest GOF out of all the chains for the two different proposal distribution strategies described in Section 5.2.1 and Section 5.2.2. We chose five combinations, instead of presenting mean and variances, to demonstrate that there is still a non-identifiability issue present.
@@ A.2 HPV model
Appendix Table 1: Calibration targets with corresponding distribution parameters ($d_{k}$ is the mean and $\sigma_{k}$ is the standard deviation). The mean and variance for each parameter were elicited from the available data independently for lack of a better alternative. For full details see section 5.1.

### v2

@@ 4 Example: SIS model
As identified by Alarid-Escudero et al. [5], the main problem in models of this type is the lack of identifiability in the parameter space, given the information available. One solution is to obtain additional data or an additional target for which data may be available. The approach adopted here is to add additional information about the parameters in the form of prior information. This prior information is specified by using information from other sources (e.g. from expert knowledge or previous experience of these parameters). Note that while the term ‘prior’ may in ordinary usage be associated with information available before collecting data (i.e. targets), in the Bayesian paradigm it is the fact that information other than contained in the targets may be incorporated that is important. The process of obtaining prior information is called prior elicitation and is a substantive area in its own right which is beyond the scope of this article. For interested readers, other literature such as a recent review[39], work on the SHELF framework [40], O’Hagan[41] and earlier articles such as Kadane and Wolfson[42] explore this area in more depth.
Consider a measles outbreak in a school setting. We aim to estimate the transmission rate ($\beta$) and recovery rate ($\gamma$) using an SIS model. Due to the limited data available from the school, the model may be susceptible to non-identifiability. To address the non-identifiability, we can incorporate prior information to constrain the parameter space. For instance for transmission rate ($\beta$) one could use expert opinion: based on previous measles outbreaks, experts might estimate that the basic reproduction number ($R_{0}$) is around 2.5. Since $R_{0}$ = $\beta$/$\gamma$, we can use this information to inform the prior distribution for $\beta$, assuming a reasonable recovery rate. Alternatively, one could look at the contact rate instead of the transmission rate itself. If we have data on the average number of contacts per student per day, we can use this to inform the prior for $\beta$, as $\beta=c\cdot p$ (where $c$ is the contact rate and $p$ is the transmission probability).
For recovery rate ($\gamma$) biological knowledge could be used. The average duration of measles infection is known, which can be used to set a prior distribution for $\gamma$ (e.g., a Beta distribution with appropriate parameters). By incorporating these prior distributions, we can help to narrow down the possible values for $\beta$ and $\gamma$, reducing the impact of non-identifiability. Sensitivity analysis can be performed to assess how the posterior distributions for $\beta$ and $\gamma$ are affected by different prior specifications.
Suppose we have a prior distribution for $\beta$ based on expert opinion and a prior distribution for $\gamma$ based on biological knowledge. By combining these priors with the likelihood function derived from the observed measles cases, we can obtain posterior distributions for $\beta$ and $\gamma$. If the posterior distribution for $\beta$ is relatively narrow, it suggests that the prior information has helped to resolve the non-identifiability issue.
One could start by using vaguely-informative uniform priors for each parameter. However, here we demonstrate how imposing an informative prior on some of the parameters improves identifiability and affects the convergence of the chains for other parameters implicitly. We impose normal priors for two parameters $c\sim N(9,1)$ and $p\sim N(0.06,0.01)$ effectively imposing a prior on $\beta$ indirectly (See right-hand side of Figure 2 for results). The chains were compared using Gelman and Rubin’s convergence diagnostic[34]. The Gelman-Rubin diagnostic (or $\hat{R}$) assesses the convergence of multiple MCMC chains by comparing between-chain and within-chain variances. If $\hat{R}$ is close to 1, it indicates convergence, meaning the chains are sampling from the same posterior distribution. Values significantly greater than 1 suggest non-convergence, requiring more iterations or adjustments. The chains without an informative additional prior imposed scored a point estimate of 2.58 with an upper confidence interval of 5.05 and the chains with normal priors scored a 1.05 point estimate with upper C.I. of 1.15.
While useful, $\hat{R}$ should be complemented by visual diagnostics like trace plots to ensure a thorough evaluation of convergence. In addition to trace plots presented in Figure 2, we also present a plot demonstrating the change from prior to posterior distributions for additional visual diagnostic. The posterior distributions for the parameters were very close to the imposed priors (Figure 3), highlighting the fact that there is little or no information in the data itself about these parameters.
Figure 3: Prior to posterior changes in distributions of 3 chains for selected parameters ($c$ and $p$) for which priors were specified directly in the SIS model calibration. All priors were normal distributions and means and standard deviations are specified on the plot. The bottom row of this figure corresponds to the right-hand panel of Figure 2.
4.1 Prior sensitivities
We conducted sensitivity analyses to assess how prior distributions influence posterior inferences. For the SIS model, we systematically varied the priors for $c$ (contact rate) and $p$ (transition probability) across a spectrum from weakly informative to strongly informative distributions. The results, as shown in Figure 3, reveal a strong dependence of posterior estimates on the prior specifications. This indicates that the available data and model structure provide insufficient information to constrain parameter estimates, leading to wide credible intervals. These findings underscore the critical importance of carefully selecting priors, especially in scenarios with limited data or high non-identifiability, to ensure meaningful and robust inferences.
@@ 4.1 Prior sensitivities
In summary, in this section, we have illustrated how the problem of non-identifiability can arise even in a relatively simple model, and shown how the use of additional information solves this problem. It is after initial runs that the modeller will detect that non-identifiability is occurring, and by carefully considering what information to add regarding other parameters, identifiability may be achieved. In practice, such additional information may come from, for example, monitoring a number of individual cases of the disease to better quantify the time to recovery. By carefully incorporating prior information, we can enhance the reliability and interpretability of disease models, even in the presence of non-identifiability. This case study demonstrates how prior knowledge can be used to guide parameter estimation and improve our understanding of disease dynamics.
@@ 5 Example: HPV model
The second model that we use in this paper to illustrate the challenges of non-identifiability in model calibration is an individual-based model, also called an agent-based model. An individual-based model simulates the characteristics of each individual and keeps track of each individual’s history, recording all information relevant to the disease in question. The advantages of agent-based models are that they can incorporate heterogeneity (through each agent), they can be physically realistic, and they allow for interventions that impact individuals differently. They are commonly used when modelling dynamic aspects of infectious diseases as described for example in Hunter et al.[43]. The challenges associated with these models are that they take more computer time to run, they can be difficult to build, and of interest to this example, they can be harder to calibrate.
The example we use here is based on the human papillomavirus (HPV) which causes cervical cancer. We follow the model developed by Kim et al. [44] which considers transitions between Healthy state, HPV infection, pre-malignant changes called cervical intra-epithelial neoplasia (CIN) 1, CIN 2 and CIN 3 (which are different levels of severity) and, finally, cancer (see Figure 4 for details). The HPV model incorporates a dynamic structure where individuals can move between different health states. Individuals can progress from a healthy state to HPV infection, and subsequently through various stages of cervical intraepithelial neoplasia (CIN) until reaching a cancer state. Importantly, the model also allows individuals to recover from HPV infection or regress from advanced CIN stages. Additionally, the model’s parameters vary by strain (i.e. transition probabilities to CIN states depend on the strain of HPV infection), and therefore the number of parameters for calibration is high. We have deliberately selected this HPV model for this paper because it is very complex and is likely to exhibit many of the challenges that were discussed in Section 2.4 and Section 2.5.
@@
Individual-based models, such as the HPV model presented here, offer a powerful framework for understanding the complex dynamics of infectious and other diseases[45, 46, 47]. By simulating the behaviour of individual agents, these models can capture heterogeneity, non-linearity, and emergent phenomena. However, the complexity of these models, often coupled with limited and population-level data, can lead to challenges in model calibration, including non-identifiability. This makes it difficult to uniquely estimate model parameters, hindering our ability to draw definitive conclusions about the underlying biological processes. Addressing these challenges is crucial for obtaining reliable model predictions and informing effective public health interventions.
5.1 Calibration using Metropolis-Hastings
In this section, we present some results of applying the Metropolis-Hastings algorithm (Appendix Algorithm 1) to the HPV model calibration. Despite the advantages of the algorithm, these results illustrate some of the problems associated with calibrating complex models using the Metropolis-Hastings algorithm described in Section 2.4 and Section 2.5.
This model has the same non-identifiability problem as the simple SIS model in Section 4. As was illustrated in Figure 1(a) of Section 2, when you have two competing transition probabilities, successfully calibrating both can be impossible if the data available for calibration is population-averaged. In our case, the only data available were the distributions of population-level observations. This is a common problem of incompatibility of population-level data (the most commonly collected and available) with agent-based models. Attempting to calibrate this model with improper uniform priors will result in chains not converging as demonstrated by the left-hand side of Figure 5. By only relying on visual diagnostics at this stage one could definitively conclude that those chains are not converging, indicating a potential problem with non-identifiability in the model. We demonstrate here how imposing even a reasonably vague prior on all parameters leads to great convergence of the chains produced by the algorithm.
Figure 5: Results of running Appendix Algorithm 1 for 120,000 steps with improper uniform priors (left-hand side) and 90,000 steps with Gamma priors (right-hand side) for the HPV model.
To address the problem of non-identifiability of the parameters as described inSection 2.5, we introduce (vaguely-)informative priors. There are many ways to incorporate any information available at a given time into priors. If the priors for each parameter are assumed to be independent, then the prior $\pi(\boldsymbol{x})$ in Appendix Algorithm 1 is just a product of one-dimensional priors $\pi(\boldsymbol{x})=\prod_{i=1}^{n}\pi_{i}(x_{i}),$ where each $\pi_{i}$ is a prior distribution for $i$’s parameter. The choice of priors represents our beliefs about parameters without taking new evidence into account. Gamma distributions were chosen to be fitted to the parameters because of their $[0,\infty)$ support and their shape. In general, Beta priors would be a standard choice for probabilities due to their $[0,1]$ support. However, this model was designed such that the baseline probabilities of transitioning were hard-coded and multipliers to those baseline probabilities were controlled by the inputs to the model, leading to this slightly unintuitive choice of prior distribution family. Priors on the model’s parameters were imposed to limit the algorithm’s search to (what we know to be) a plausible region of the space. The uniqueness of such a region can depend on many factors. Possible sources of information from which one could construct priors also include expert opinions and alternative data sets.
We specify informative prior distributions for transition rates in our HPV model using the study by Jackson et. al [48]. This paper demonstrated the efficacy of using Bayesian methods to combine diverse evidence and propagate uncertainties in model calibration. Building on their results, we used Gamma distributions for prior specifications consistent with the identified intervals, considering their suitability for non-negative values and their flexibility in capturing various levels of uncertainty. By carefully selecting Gamma parameters based on expert knowledge and available data, we were able to incorporate valuable prior information into our model, enhancing the reliability of our parameter estimates and improving the overall robustness of our HPV model calibration. As highlighted above, the prior distribution can be thought of as ‘other information’ about the parameters than that provided by the likelihood. In this case, structural priors can be placed upon the parameter space by observing that large sections are not compatible with plausible data values. Plausible values had also to be chosen for the backward transitions.
The posterior distribution of the chains of the Metropolis-Hastings algorithm along with the corresponding priors for four selected parameters can be found in Figure 6. Although here we focus on four of the most interesting parameters, the resulting posterior densities for all 26 parameters are presented in Appendix Figure 1. It is clear from Figure 6 that all the posterior distributions are distinct from the priors and in some cases, despite the somewhat vague priors, very distinct. This highlights again the point made in Section 4 and Figure 2. In a complex model like this, priors of interacting parameters could have a knock-on effect. Even though there was not enough information in the data and the model itself to identify good parameter values, and the prior on this parameter was vague, there was enough information in other parameter’s priors to have a knock-on effect and for the posterior distribution to have a pronounced peak, especially for CIN23 to Local Cancer (HR 16).
It is also visible from the density plots in Figure 6 (and even more so in Appendix Figure 1) that some multi-modality exists in the posterior distributions. All three chains recognise the multi-modality and sampling from the same distribution. This multi-modality was not picked up by the original chain nor was it present in the Gamma priors. This highlights the importance of prior elicitation in complex models. The balance between informing some parameters from the data or informing them through the priors could be drastically shifted towards the priors in a situation of non-identifiability[5] or just general lack of information in the available data.
Figure 6: Prior to posterior changes in distributions of 3 chains for selected 4 parameters for the HPV model. All priors were Gamma distributions. Posterior distribution plots for all parameters can be found in Appendix Figure 1.
As a small aside we would like to make a point here that despite the complications associated with direct application of the Metropolis-Hastings algorithm (with improper uniform priors) to the HPV model calibration, compared to the original method of Kim et al. [44] the goodness-of-fit (GOF) measure (see Appendix B for details) was consistently better for all chains (low scores being better), staying near 20, fluctuating up and down around that value. Despite the problems, this was true for all approaches, even the ones that we have chosen not to present here. If we compare this to the simple Monte Carlo approach of Kim et al. [44], the best fitting parameter values had GOF values around 122 for the Irish data (with minimal value at 72 and maximal value at 185 in a sample of 133 best fitting values). The corresponding p-values are 93.58% and $\ll 0.01\%$ for our approach (GOF=20) and Kim et al. [44] approach (GOF=122) respectively. However, the fit of our approach appears to be almost ‘too good’ relative to $\chi^{2}_{31}$. That can be explained by the correlation in the targets, so if one is fitting well, a few others are also fitting well. If we remove the assumption of independence, that measure of GOF would have a different distribution (depending on the correlations), but the mean of that distribution will be shifted towards zero compared to $E(\chi^{2}_{k=31})=31$. This will make the value of GOF=20 more realistic.
By incorporating informative priors, we were able to effectively address the challenges of non-identifiability and improve the precision of our parameter estimates. We used Gamma distributions to specify priors for transition-rate multipliers, leveraging expert knowledge and available data. These priors helped to constrain the parameter space and prevent the algorithm from exploring implausible regions. As a result, the posterior distributions were more focused and informative, leading to more reliable model predictions and inferences for decision-makers.
6 Discussion
The Bayesian approach to model calibration is becoming more popular[49] and more research is focusing on MCMC techniques in particular. Several calibration approaches become more difficult to apply in the context of stochastic models, such as stochastic microsimulations or agent-based models[49]. Given the advantages of a Bayesian stochastic perspective, and in particular, here an MCMC approach, the fact that formal convergence is not always easily obtained without substantial additional work requires much attention. Such work includes better data collection or data representation in terms of priors as well as tuning the MCMC to the problem.
One of the main observations in this work was that it is possible to explicitly incorporate additional information into the inference in the form of priors during the calibration process and that this had a significant effect on the algorithm’s performance. This has a big impact on the issue of non-identifiability. Once the additional information was included (here through the revision of the prior), the algorithm was not impaired by the presence of some local multi-modality and all three chains sampled from the same distribution. This led to a desirable effect of convergence that was coupled with the improved acceptance rate of the algorithm and a visible lack of autocorrelation in the sample.
The fact that the application of the Metropolis-Hastings algorithm yielded results with a lower GOF than the simple Monte Carlo approach is not surprising, given that the Metropolis-Hastings algorithm is designed to lead the sampler to a high posterior probability area, whether it is a global or a local maximum. Additionally, this technique provides probabilistic estimates of the parameters of interest and reflects their joint uncertainty. Usually, the acceptance criterion is applied after the fact, which leads to inefficiency.
In a population-level model of Hepatitis C infection, a similar process of calibration was performed [50]. This case had none of the problems that we had with the HPV model calibration. This was mostly because good priors were provided by an expert. Furthermore, because the Hepatitis C model did not have any backward transitions, there were fewer problems with identifiability, compared to the HPV model. Another possible advantage of that model was its relative simplicity, making it difficult to over-parameterise, compared to the individual-based models. This allowed us to try multiple strategies, before selecting the best one. We hypothesise that the relatively simpler structure of the Hepatitis C model, compared with the HPV model, yields a smoother probabilistic link between the model parameters and outputs. Of course, model choice is very influential in model calibration and different model types might yield different calibration results.
Funding
Acknowledgements
We would like to thank MACSI (University of Limerick) and the Health Research Board who have provided the funding for this research. We are also grateful to Jane Kim for sharing the code from their paper[44]. Daria Semochkina would also like to thank Paul A. Smith (for his advice on article preparation and many helpful discussions) and John G. Donohue (for tirelessly proofreading this article countless times). We would also like to thank the Editor-in-chief and anonymous reviewers for their helpful contribution to shaping this paper.
Appendix A Metropolis-Hastings algorithm
The central idea of Algorithm 1 is to locally update, i.e. use the accepted parameter value $\boldsymbol{x}_{(t)}$ at time step $t$ to generate the next proposed value. This makes it easier to think of a suitable proposal step (that is, of course, conditional), albeit with the disadvantage of producing a Markov chain (as opposed to an independent realisation). This Markov chain converges to the posterior distribution of interest after enough steps. The sample values are taken from the Markov chain after convergence. The basic steps are presented in Algorithm 1.
Algorithm 1 Metropolis-Hastings algorithm
Starting from $\boldsymbol{x}_{(0)}=(x_{(0)}^{1},x_{(0)}^{2},...,x_{(0)}^{n}),$ for $t=1$ to $N$ do
1:Draw $\boldsymbol{x}\sim q(\cdot|\boldsymbol{x}_{(t-1)})$
2:Compute
$p(\boldsymbol{x}|\boldsymbol{x}_{(t-1)})=\min\left\{1,\frac{f(\boldsymbol{x})q(\boldsymbol{x}_{(t-1)}|\boldsymbol{x})}{f(\boldsymbol{x}_{(t-1)})q(\boldsymbol{x}|\boldsymbol{x}_{(t-1)})}\right\}$ (1)
3:With probability $p(\boldsymbol{x}|\boldsymbol{x}_{(t-1)})$ accept the step and set $\boldsymbol{x}_{(t)}=\boldsymbol{x}$, otherwise reject the step and let $\boldsymbol{x}_{(t)}=\boldsymbol{x}_{(t-1)}$
End
Appendix B HPV model
@@
Here, we present further details about the HPV model, described in Section 5. We assume that all of the targets have normal distributions $f_{k}\sim N(d_{k},\sigma_{k}),$ centred at the observed data $d_{k}$ and with variances $\sigma_{k}$ that were calculated from the data. We assume all targets to be independent. Because of that assumption, our likelihood is just a product of individual likelihoods
@@
The details of the means and standard deviations can be found in Appendix Table 1. If we simplify the GOF measure, we can see that it is $\chi^{2}$ with 31 degrees of freedom (assuming independence).
@@
Appendix Table 1: Calibration targets with corresponding distribution parameters ($d_{k}$ is the mean and $\sigma_{k}$ is the standard deviation). The mean and variance for each parameter were elicited from the available data independently for lack of a better alternative.